# Towards Fair In-Context Learning with Tabular Foundation Models

### v2

@@ 3.3 In-context prediction
After performing demonstration selection or transformation using the fairness intervention methods presented previously, we pass them through the tabular foundation model as in-context examples for predicting class labels on the test set. In this paper, we consider TabPFN (Hollmann et al., 2025) and TabICL (Qu et al., 2025) as tabular foundation models. These models use Transformer architectures as their backbones and are trained with extensive synthetic datasets generated from structured causal networks. While TabPFN can handle a maximum of 10k samples with 500 features, TabICL can handle up to 500K samples.
@@ 4.1 Experimental Setup
More details about each dataset, including the sensitive attributes used for fairness evaluation, the number of samples, and the number of features, can be found in the Appendix in Table 1.
Metrics
In addition to the classic accuracy, we consider three group fairness metrics, i.e., Demographic Parity (DP), Equal Opportunity (EOP), and Equalized Odds (EOD), more details can be found in B.1.
Evaluation
For evaluation, we split each dataset into 80%-20%, where the 20% is held out and used to train the sensitive attribute classifier for uncertainty quantification. For the remaining 80% we use 5-fold cross-validation with random shuffle across ten independent runs. This ensures our evaluation is robust and reliable since every data point is used as a test or in-context example across k-fold evaluations. We report the average and standard deviation of fairness and accuracy performance across the k-fold test sets and the ten random independent seeds. As aforementioned, we use TabPFN and TabICL as foundation models for ICL prediction.
Baselines
Uncertain: using the uncertainty of the sensitive attribute prediction to select in-context examples. We use the Mappie implementation of conformal prediction (Cordier et al., 2023) to estimate the uncertainty of the training data. We fix the coverage parameter , of conformal prediction to and only select as in-context example samples with prediction sets equal to two, i.e., samples with uncertain sensitive attributes. We consider two variants of the method under different model classes used to train the sensitive attribute classifier for uncertainty estimation: a variant that uses the traditional Logistic Regression model (Uncertain+LR) and a variant that uses a foundation model (Uncertain+TabPFN).
4.2 Results and discussion
We evaluate several aspects of fairness in ICL prediction with foundational tabular datasets. First, we compare the different baselines considered in terms of fairness and accuracy; For the methods with controllable tradeoff between fairness and accuracy, we vary the hyperparameter controlling the fairness-accuracy tradeoff and compare their Pareto fronts. We then compare the foundation model under the best-performing fairness intervention method. Finally, we provide an ablation study on the impact of the size of the in-context set on fairness and accuracy.
4.2.1 Baseline comparison
Figure 1 summarizes the accuracy and fairness of the different baselines considered across the eight datasets with TabPFN as the foundational model. The group-balanced method shows a marginal improvement in fairness compared to the Vanilla method across datasets. This shows that representation bias is not the only source of bias in the data, and having a group-balanced in-context sample set does not necessarily reduce performance disparity across groups. The unfairness remaining could be due to other sources of bias such as historical bias and measurement bias (Mehrabi et al., 2022).
Surprisingly, Correlation Remover exacerbates unfairness in most datasets. For example, on the ACSIncome dataset, demographic parity increases from to . To further investigate the failure case of Correlation Remover, we measure and compare the reconstruction of sensitive attributes by the foundation model before and after applying the preprocessing intervention. Specifically, we perform ICL using the sensitive attribute as the target variable and use the accuracy of the ICL prediction of the sensitive attribute in the test set as a measure of how the foundation model can reconstruct the sensitive attribute after the fairness intervention. We observe that TabPFN and TabICL can fully reconstruct (up to accuracy) the sensitive attribute based on the feature transformation applied by Correlation Remover (see Table 2 in Appendix). This demonstrates that transformer-based models like TabPFN, pretrained on synthetic data with different sample and feature interactions, can uncover the transformation applied to the non-sensitive features at test-time. By transforming every non-sensitive feature based on the sensitive one, Correlation Remover introduces a sensitive information leakage that the foundational model uncovers and relies upon for predicting the target variable, thereby exacerbating unfairness. Further discussion and results can be found in Appendix D, where we also demonstrate that applying the feature transformation only to the training set and leaving the test set unchanged yields better fairness outcomes compared to the Vanilla approach.
In Appendix C we provide an ablation on the size of in-context examples showing that our findings remain consistent across smaller in-context sizes. More specifically, accuracy significantly increases with the size of the in-context examples, while fairness steadily increases and becomes almost constant when the in-context size exceeds 700 examples.
4.2.2 Fairness Accuracy Tradeoff
In the previous experiment, we fixed the parameters and , controlling the tradeoff between fairness and accuracy for the Correlation Remover and the Uncertain methods, respectively. For a better comparison of the fairness accuracy performances, we consider a range of values between [0, 1] for parameters controlling the tradeoff and plot the Pareto front. Figure 2(a) and 2(b) show the Pareto front on the ACSIncome and ACSMobility datasets, respectively, and compare the Vanilla ICL method, Correlation Remover, and the two variants of the Uncertain method. Results on other datasets can be found in the Appendix (Figure 6) and results using TabICL in Figures 9 and 8. Each scatter point in the figure represents a different value of and for the corresponding ICL method. As can be seen, both variants of the Uncertain consistently have better Pareto dominant points while Correlation Remover can perform worse than the Vanilla ICL in terms of fairness, even with different values of . This further provides evidence of bias amplification of Correlation Remover when used with foundation models. We also see that the Uncertain method can consistently control the fairness accuracy trade-off with smaller values of , enforcing better fairness at the expense of accuracy. The slight decrease in accuracy when the Uncertain method is applied is mainly due to the reduced size of the in-context set, since more and more samples with higher certainty of sensitive attributes are not included in the in-context set.
4.2.3 Foundation Model Comparison: TabICL vs. TabPFN
In the previous experiment, we compared Correlation Remover and Uncertain methods in terms of fairness-accuracy tradeoff by varying their parameter controlling the tradeoff. The results demonstrated that the Uncertain methods often provide better Pareto dominant points compared to the Vanilla ICL and Correlation Remover. And the Uncertain using TabPFN for uncertainty measurement (Uncertain+TabPFN) tends to perform better than using Logistic Regression for uncertainty estimation (Uncertain+LR). In this experiment, we apply the same setup and compare the tradeoffs for the different foundation tabular models under the best-performing fairness intervention, i.e., Uncertain+TabPFN. Figure 3(a) and 3(b) show the fairness accuracy Pareto front on the ACSIncome and ACSMobility datasets, respectively.
We can observe that both foundation models tend to have similar fairness performances while TabPFN often provides better accuracy across datasets. This is consistent with the observation made by the TabICL’s authors when comparing with the TabPFN (Qu et al., 2025). We observe similar trends on other datasets, and the results can be found in Figure 10 of the Appendix.
5 Conclusion and future works
Acknowledgement
Ethics Statement
Appendix
Appendix A Datasets
@@ B.2 Correlation Remover
is the mean vector of the sensitive features.
@@
Appendix C Ablation on impact in-context sample size
All the previous experiments used the entire training as in-context examples when possible. Note that the current version of TabPFN cannot handle more than samples (Hollmann et al., 2025). As an ablation study, we vary the in-context sample size ([100, 300, 500, 700, 1500, 2000, 2500, 3000, 4000, 5000]) and measure fairness and accuracy using the same experimental setup as in Section 4.2.1. Figure 9 shows that accuracy significantly increases when more samples are added to the in-context set. Unfainness increases slightly before remaining almost constant when the in-context set size exceeds . We observe that Correlation Remover still has the highest unfairness across in-context size and data. On the other hand, Uncertain+TabPFN method consistently has the lowest fairness violation across in-context sizes, demonstrating its robustness to the size of training data points.
Appendix D On the failure of Correlation Remover
In the main paper, we observed that the transformation 10 on both the training and the testing data can exacerbate unfairness in ICL predictions. We hypothesized that the foundation model inferred the sensitive attribute from the linear transformation applied to each non-sensitive feature, resulting in higher unfairness in the ICL predictions. To validate this hypothesis, we perform ICL prediction of the sensitive attribute after the fairness interventions are applied. As can be seen in Table 2, ICL prediction of the sensitive attribute results in 100% accuracy after Correlation Remover is applied. These results suggest the foundation model still relies on the sensitive attribute after correlation removal is performed.
@@ Appendix D On the failure of Correlation Remover
| Dataset | Correlation R. | TabPFN | TabICL | ||
| Accuracy | F1 Score | Accuracy | F1 Score | ||
| ACSIncome | None | ||||
| Train & Test | |||||
| Train | |||||
| ACSTravelTime | None | ||||
| Train & Test | |||||
| Train | |||||
| ACSPublicCoverage | None | ||||
| Train & Test | |||||
| Train | |||||
| ACSEmployment | None | ||||
| Train & Test | |||||
| Train | |||||
| ACSMobility | None | ||||
| Train & Test | |||||
| Train | |||||
| Diabetes | None | ||||
| Train & Test | |||||
| Train | |||||
| German Credit | None | ||||
| Train & Test | |||||
| Train | |||||
| CelebA | None | ||||
| Train & Test | |||||
| Train | |||||
For further verification, we consider a variant of Correlation Remover where we apply the feature transformation (Eq. 10) only to the training dataset, leaving the test data unchanged. Table 3 shows this variant (Eq. 10) significantly reduces the accuracy of ICL predictions. This demonstrates that the foundation model uses the transformation applied to the testing set as a proxy to fully reconstruct the sensitive attribute. Aïvodji et al. (2021) discuss several scenarios where one can apply a data transformation to reduce correlation with the sensitive attribute. More specifically, they considered scenarios where the transformation is applied (1) both to the training and testing dataset, (2) only to the training set, and (3) only to the test set. We evaluated the fairness accuracy tradeoff of applying feature transformation only to the training set. The results in Figure 5 show that this variant can improve fairness compared to applying the transformation to both the training and testing datasets. This further illustrates the ability of the foundation models to reconstruct the sensitive attributes when the transformation is applied to the test set.
Appendix E Supplementary resutls

### v3

@@ 3.3 In-context Prediction
After performing demonstration selection or transformation using the fairness intervention methods presented previously, we pass them through the tabular foundation model as in-context examples for predicting class labels on the test set. In this paper, we consider three tabular foundation models: TabDPT (Ma et al., 2024), TabICL (Qu et al., 2025), and TabPFN v2 (Hollmann et al., 2025). While all these models use Transformer architectures as their backbones and are pretrained on a large amount of datasets, TabICL and TabPFN use only synthetic data generated from structured causal networks, and TabDPT uses real-world data. The current version of TabPFN can handle a maximum of 10k samples with 500 features, and TabICL can handle up to 500K samples. We randomly subsample in-context examples when the context set size exceeds the maximum number of samples allowed by the foundation model being evaluated.
@@ 4.1 Experimental Setup
We experiment on tasks from the recently proposed folktables (Ding et al., 2022), which contains data extracted from the American Community Survey (ACS) Public Use Microdata Sample (PUMS) (Ding et al., 2022). More specifically, we experiment with the following ACS PUMS tasks: ACSIncome, ACSMobility, ACSTravelTime, ACSEmployment, ACSPublicCoverage. These tasks reflect a range of real-world predictive challenges with fairness concerns. A limitation of the ACS PUMS datasets is that they are US-centric; we diversify the experimental setup by including other tasks and datasets. Specifically, we also experiment on the other tabular datasets and tasks including: Diabetes (Gardner et al., 2023), German Credit (Frank, 2010), and CelebA (Liu et al., 2018). More details about each dataset, including the sensitive attributes used for fairness evaluation, the number of samples, and the number of features, can be found in the Appendix A.
Metrics
In addition to the classic accuracy, we consider three group fairness metrics, i.e., Demographic Parity (DP), Equal Opportunity (EOP), and Equalized Odds (EOD). More details about group fairness metrics can be found in B.1.
Evaluation
For evaluation, we split each dataset into 80%-20%, where the 20% is held out and used to train the sensitive attribute classifier for uncertainty quantification. For the remaining 80% we use 5-fold cross-validation with random shuffle across ten independent runs. This ensures our evaluation is robust and reliable since every data point is used as a test or in-context example across k-fold evaluations. We report the average and standard deviation of fairness and accuracy performance across the k-fold test sets and the ten random independent seeds. As aforementioned, we use TabICL, TabDPT and TabPFN as foundation models with their respective default parameters.
Baselines
Uncertain: using the uncertainty of the sensitive attribute prediction to select in-context examples. We use the Mapie (Taquet et al., 2022) implementation of conformal prediction (Cordier et al., 2023) to estimate the uncertainty of the training data. We fix the coverage parameter , of conformal prediction to and only select as in-context example samples with prediction sets equal to two, i.e., samples with uncertain sensitive attributes. We consider two variants of the method under different model classes used to train the sensitive attribute classifier for uncertainty estimation: a variant that uses the traditional Logistic Regression model (Uncertain+LR) and a variant that uses a foundation model (Uncertain+TabPFN).
4.2 Results and Discussion
We evaluate several aspects of fairness in ICL prediction with foundational tabular datasets. First, we compare the different baselines considered in terms of fairness and accuracy; For the methods with controllable tradeoff between fairness and accuracy, we vary the hyperparameter controlling the fairness-accuracy tradeoff and compare their Pareto fronts. We then compare the foundation model under the best-performing fairness intervention method. Finally, we provide an ablation study on the impact of the size of the in-context set on fairness and accuracy and discuss the failure case of the correlation removal method.
4.2.1 Baseline Comparison
Figure 1 summarizes the accuracy and fairness outcomes of various baselines across eight datasets, using TabPFN as the foundation model. The group-balanced method yields only marginal improvements in fairness compared to the Vanilla ICL approach, suggesting that representation bias alone does not account for the observed disparities. In other words, even when the in-context sample set is group-balanced, performance disparities across subgroups persist—likely due to other sources of bias such as historical or measurement bias (Mehrabi et al., 2022).
More strikingly, the Correlation Remover method often exacerbates unfairness. For instance, on the ACSIncome dataset, demographic parity disparity increases from to following the application of Correlation Remover. To investigate this failure mode, we assess whether the foundation model can reconstruct the sensitive attribute after Correlation Remover ’s preprocessing step. Specifically, we perform ICL where the sensitive attribute is treated as the prediction target, and measure the model’s test accuracy in reconstructing it.
We find that TabPFN and TabICL can reconstruct the sensitive attribute with up to accuracy (see Table 4 in Appendix), even after Correlation Remover ’s intervention. This suggests that these transformer-based models, pretrained on synthetic data with diverse structural priors, are capable of uncovering and leveraging the transformation patterns introduced by Correlation Remover. Since Correlation Remover modifies each non-sensitive feature based on the sensitive attribute, it unintentionally leaks sensitive information. The foundation model learns to exploit this leakage, ultimately relying on the hidden sensitive signal to predict the target variable—thereby amplifying unfairness.
Further results and discussion are presented in Section 4.5, where we show that applying the feature transformation only to the training set—while keeping the test set unchanged—leads to improved fairness outcomes compared to the Vanilla baseline.
4.2.2 Fairness-Accuracy Tradeoff
In the previous experiment, we fixed the parameters and , which control the tradeoff between fairness and accuracy for the Correlation Remover and Uncertain methods, respectively. To better evaluate this tradeoff, we now vary and over the interval and compare the resulting Pareto fronts of each method.
Using TabPFN as the foundation model, Figures 2(a) and 2(b) display the Pareto fronts on the ACSIncome and ACSMobility datasets, respectively. These figures compare the Vanilla ICL method, Correlation Remover, and the two variants of our uncertainty-based method: Uncertain+LR and Uncertain+TabPFN. Each scatter point corresponds to a different value of or for a given method.
Both variants of the Uncertain method consistently achieve Pareto-dominant points compared to alternatives. In contrast, Correlation Remover can underperform even the Vanilla ICL in terms of fairness, suggesting that it may amplify bias when used with foundation models. Notably, our uncertainty-based method provides a more stable control over the fairness–accuracy tradeoff across values, enabling improved fairness with relatively modest reductions in accuracy.
This accuracy drop is primarily attributed to the smaller in-context set size: as decreases, more samples with high sensitive attribute certainty are excluded from the context, effectively reducing the number of training examples. As discussed further in Section 4.3, the size of the in-context set has a strong influence on model accuracy.
Among the two Uncertain variants, Uncertain+TabPFN —which uses TabPFN for uncertainty estimation—tends to yield better Pareto fronts. This suggests that TabPFN functions as a stronger conformal predictor than logistic regression, allowing for more accurate estimation of sensitive attribute uncertainty and thus more effective in-context selection for fairness.
4.2.3 Comparison of Foundation Models: TabICL vs. TabDPT vs. TabPFN
In the previous experiment, we compared the Correlation Remover and Uncertain methods in terms of their fairness–accuracy tradeoffs by varying their respective control parameters. The results showed that the Uncertain methods often achieved better Pareto-dominant points compared to both Vanilla ICL and Correlation Remover. Notably, using TabPFN for uncertainty estimation (Uncertain+TabPFN) outperformed the variant that relies on logistic regression (Uncertain+LR).
In this experiment, we extend the comparison to different foundation models by applying the same setup using our best-performing fairness intervention: Uncertain+TabPFN. Figures 3(a) and 3(b) show the fairness–accuracy Pareto fronts on the ACSIncome and ACSMobility datasets, respectively.
Without any fairness intervention (i.e., using the vanilla ICL approach), there is no clear winner among the foundation models in terms of fairness; the results vary across datasets. While one might expect TabDPT —pretrained on real-world data—to encode more real-world biases and thus exhibit poorer fairness performance, this is not consistently observed. In fact, TabDPT often performs comparably to or even better than models pretrained on synthetic data.
Under fairness intervention with Uncertain+TabPFN, all three foundation models exhibit similar fairness peformance, demonstrating that the Uncertain method can consistently control the fairness–accuracy trade-off regardless of the underlying foundation model. However, both TabPFN and TabDPT consistently achieve higher accuracy across datasets. These findings are consistent with previous studies: the authors of TabICL reported strong performance for TabPFN (Qu et al., 2025), while the authors of TabDPT highlighted its competitive performance relative to TabPFN.
We observe similar patterns across additional datasets and results are provided in the Appendix (see Figure 6).
4.3 Ablation on Impact In-context Sample Size
In all previous experiments, we used the full training set as the in-context example set whenever possible. For example, the current version of TabPFN is limited to handling a maximum of 10,000 samples (Hollmann et al., 2025). To understand the impact of in-context set size, we conduct an ablation study by varying the number of in-context examples across the range [100,300,500,700,1500,2000,2500,3000,4000,5000], while keeping the evaluation setup identical to that described in Section 4.2.1.
Figure 4 shows that accuracy increases substantially with larger in-context sets, particularly in the lower range. However, unfairness shows a slight increase initially and then stabilizes once the in-context set exceeds approximately 700 examples. Across all in-context sizes and datasets, the Correlation Remover method consistently exhibits the highest level of unfairness, reinforcing earlier observations of its bias amplification. In contrast, the Uncertain+TabPFN method consistently achieves the lowest fairness violation, demonstrating robustness to the size of the in-context set. This indicates that even when only a subset of training data can be used, Uncertain+TabPFN remains effective at mitigating bias without overly compromising accuracy.
4.4 Comparison with LLM-based ICL Methods
Despite LLMs’ in-context learning capabilities, they often lag behind classical machine learning models on large-scale tabular datasets (Hegselmann et al., 2023). To validate the effectiveness of our uncertainty-based demonstration selection method, we compare it against fairness-aware LLM-based fair ICL approaches proposed by Hu et al. (2024) and Bhaila et al. (2024) on the widely studied Adult dataset (Asuncion et al., 2007). Table 1 presents the results. Vanilla TabPFN achieves an accuracy of , significantly outperforming the LLaMA-2-13B baseline, which reaches only . This 10-point gap underscores the performance advantage of specialized models for tabular data, consistent with prior findings that LLMs struggle on such inputs when used naively (Hegselmann et al., 2023).
While these comparisons rely on LLMs that are not the most recent generation, and performance may have improved with newer models, the results nonetheless highlight the continued strength of tabular foundation models like TabPFN in this setting. Moreover, when fairness interventions are applied, our uncertainty-based selection strategy consistently preserves high predictive performance while achieving comparable or better fairness outcomes than the LLM-based fair ICL approaches—demonstrating that principled sample selection can improve fairness without sacrificing accuracy.
4.5 On the Failure of Correlation Remover
| Dataset | Fairness Intervention | TabPFN | TabICL | ||
| Accuracy | F1 Score | Accuracy | F1 Score | ||
| ACSIncome | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
| ACSTravelTime | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
| ACSPublicCoverage | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
| ACSEmployment | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
| ACSMobility | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
| Diabetes | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
| German Credit | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
| CelebA | None | ||||
| Correlation R. (S1) | |||||
| Correlation R. (S2) | |||||
In our previous experiments, we observed that applying the correlation removal (CR) to both the training and test data can exacerbate unfairness in ICL predictions. We hypothesized that the foundation model infers the sensitive attribute from the linear transformation applied to each non-sensitive feature, which inadvertently leaks sensitive information and increases unfairness. To validate this hypothesis, we conducted correlation removal prediction of the sensitive attribute after applying fairness interventions. As shown in Table 4, the ICL prediction of the sensitive attribute achieves accuracy following the application of CR. This indicates that the foundation model continues to rely heavily on the sensitive attribute even after correlation removal is performed.
For further verification, we tested a variant of correlation removal where the feature transformation (see Eq. 10 in Appendix) is applied only to the training data, leaving the test data unchanged (referred to as variant S2). Table 2 shows that this variant significantly reduces the ICL prediction accuracy of the sensitive attribute. This demonstrates that the foundation model exploits the transformation applied to the test set in the original correlation removal method (variant S1) as a proxy to fully reconstruct sensitive attributes.
This observation aligns with prior discussions by A¨ıvodji et al. (2021) regarding scenarios where data transformations reduce correlation with sensitive attributes. Besides the classical correlation removal setting (S1), where transformations are applied to both training and test sets, we evaluated the fairness–accuracy trade-off when applying the transformation only to the training set (S2).
As illustrated in Figure 5, variant S2 of CR substantially improves fairness compared to S1. This confirms the foundation models’ capacity to reconstruct sensitive information when the test set is transformed, leading to bias amplification. And the drop in reconstruction accuracy of the sensitive attribute directly supports the observed improvement in fairness when using variant S2
Based on these findings, we recommend practitioners apply the S2 variant of Correlation Remover—transforming only the training data—when using correlation removal in in-context learning frameworks, to avoid sensitive attribute leakage and mitigate bias amplification during testing.
5 Conclusion and future works
Acknowledgement
Ethics Statement
Appendix
Appendix A Datasets
@@ B.2 Correlation Remover
is the mean vector of the sensitive features, i.e., is the mean of -th the sensitive feature.
@@
B.3 Uncertainty measure with conformal prediction
While any uncertainty measurement method could be used in our method, we employ conformal prediction due to its strong coverage guarantees. Specifically, we used Split Conformal Prediction (Angelopoulos et al., 2023), which is a distribution-free and model-agnostic method that provides prediction sets for classification tasks, ensuring a user-specified coverage level with no assumptions beyond data exchangeability.
Given labeled data with binary sensitive attribute we split the data into proper training set ( ) and calibration set (). In the experiments, we did a 50%-50% split of the dataset with sensitive attribute to obtain and .
We then train a probabilistic classifier , on , yielding predictions:
In our setup, could be Logistic Regression (Uncertain+LR) or TabPFN (Uncertain+TabPFN). The starting point for conformal prediction is what is called a nonconformity measure, a real-valued function that measures how a prediction is different from any possible class label.
Nonconformity Scores
After training on the proper training set, we use the calibration dataset to compute the nonconformity scores, which measure how far the prediction is from the true label. More specifically, for each calibration point , we considered the nonconformity score is defined as:
We then compute the quantile threshold based on the user-defined target coverage , e.g., mean 95% coverage. Specifically, is defined based on quantile of the nonconformity scores.
Prediction Set
The quantile threshold is used to build the prediction set of data points from the test set. More specifically, for a test sample , we compute the prediction probability and derive its prediction set as follows:
When the prediction set only contains {0} or {1}, then the prediction is confident with at least probability, while when the prediction set contains both labels, i.e., {0, 1}, the prediction is considered uncertain. Our uncertainty-based demonstration selection method uses only samples whose prediction set contains two values. The coverage guarantee of conformal prediction holds under the assumption that the calibration and test data are exchangeable. We use the open-source implementation of split conformal prediction provided by the MAPIE Python package.
Example.
Consider a simplified task consisting of two non-sensitive ( and ) features one binary target () and a binary sensitive attribute ().
We first train a sensitive attribute classifier, using a fraction of the dataset (20% in our experiment), to estimate the uncertainties. We train the sensitive attribute classifier using and to predict .
A conformal predictor (details provided above) will return for a given (test) sample a prediction set that contains the true sensitive attribute with probability of at least . For example, setting means 95% of the data will contain the true label in their prediction sets. The size of the prediction set, therefore, provides information about the prediction uncertainty, i.e., means the prediction is confident with at least probability, and means the prediction is uncertain. The intuition of our method is that using samples with uncertain predictions makes it challenging for the foundation model to infer and rely on the sensitive attributes to make predictions on the target.
We therefore filter out a demonstration example if and select it if . helps to control the fairness-accuracy tradeoff since smaller corresponds to dropping more demonstration examples with higher confident sensitive attributes prediction.
Appendix C Supplementary resutls
@@
The results in Figure 1 and Table 5 show that improving fairness in the ACSPublicCoverage and Diabetes datasets is particularly challenging. We observe that unfairness is already pretty small, and the fairness methods alter the training data, either by transforming the features (correlation removal) or reducing its size (group-balanced and Uncertain methods), which reduces accuracy, without necessarily improving on fairness. For Table 1, we used a fixed value for the hyperparameter controlling the fairness accuracy tradeoff, for the uncertain method and for correlation removal. Figure 7(c) and 7(d) show that for different values of , there are Pareto-optimality points of the uncertain method that significantly reduce unfairness compared to random demonstration selection baselines. These points correspond to values of . In sum, for some tasks, in a low unfairness regime such as PublicCoverage and Diabetes, smaller values of are needed to reduce unfairness, which means more data points with a confident prediction of the sensitive attribute must be removed from the demonstration set. The value used in practice can be defined by stakeholders, decision-makers or business needs, considering that smaller means increased fairness and a drop in accuracy, and higher means a decrease in fairness and higher accuracy.